### 20 Apr 2026 v2

# Wickstead’s conjecture on positive projections and non-representable Banach lattice algebras

Abstract.
Let $X$ be a Dedekind complete Banach lattice, and let $P\colon X\to X$ be a positive projection for which the largest central operator below $P$ is $\alpha\operatorname{id}_{X}$, for some $\alpha\geq 0$. Wickstead conjectured that $\alpha$ must either be 0 or $1/n$, for some $n\in\mathbb{N}$, and proved it for finite-dimensional $X$. In this paper, we show that the conjecture holds in general. As a consequence, we settle the representation problem for Banach lattice algebras: we show that there exist Banach lattice algebras of dimension $2$ that do not admit a faithful representation as regular operators on any Dedekind complete Banach lattice.
Key words and phrases:
positive operator, positive projection, Banach lattice, Banach lattice algebra
2020 Mathematics Subject Classification:
47B65, 46B42, 46A40, 06F25
Research supported by an FPI–UAM 2023 contract (funded by Universidad Autónoma de Madrid) and by grants PID2024-162214NB-I00 and CEX2023-001347-S (funded by MCIN/AEI/10.13039/501100011033).
1. Introduction
In [undefj], A. W. Wickstead formulated the following question. Let $X$ be a Dedekind complete Banach lattice, and let $P\colon X\to X$ be a positive projection for which there exists an $\alpha\geq 0$ satisfying:
  1. (i)
    $\alpha\operatorname{id}_{X}\leq P$, and
  2. (ii)
    whenever $T$ is a positive operator such that both $T\leq\lambda\operatorname{id}_{X}$, for some $\lambda\geq 0$, and $T\leq P$, then also $T\leq\alpha\operatorname{id}_{X}$.
Which values can $\alpha$ take? Wickstead conjectured that either $\alpha=0$ or $\alpha=1/n$, for some $n\in\mathbb{N}$. The first main goal of this paper is to prove that this is indeed the case (3.1).
When $X$ is finite-dimensional, conditions (i) and (ii) imply that the diagonal elements in the matrix representing $P$ are all equal to $\alpha$. Thus we will refer to these two conditions by saying that $P$ has constant diagonal $\alpha$. When $X=C(K)$, for some compact Hausdorff space $K$, or $X=L_{1}(\mu)$, for some finite measure $\mu$, that $P$ has constant diagonal $\alpha$ means that the biggest multiplication operator (by an element of $C(K)$ or of $L_{\infty}(\mu)$, respectively) below $P$ is $\alpha\operatorname{id}_{X}$.
The motivation behind this question lies in the representation of Banach lattice algebras. A Banach lattice algebra is a Banach lattice together with a Banach algebra structure for which the product of positive elements is positive. One of the most prominent questions regarding Banach lattice algebras is whether every such object can be faithfully represented (both as a lattice and as an algebra) in $\mathcal{L}_{r}(X)$, the space of regular operators on some Dedekind complete Banach lattice $X$ [undefh]. This pursuit of concrete realization is heavily inspired by fundational representation theorems in operator algebras, where abstractly defined structures are successfully shown to be isomorphic to concrete algebras of operators. A prime example is the Gelfand–Naimark–Segal theorem, which establishes that any abstract $C^{*}$-algebra can be faithfully represented as a closed subalgebra of bounded operators on a Hilbert space. A positive answer to the representation problem for Banach lattice algebras would provide a similar structural cornerstone, allowing abstract lattice algebras to be analyzed through the concrete, well-understood properties of regular operators on Dedekind complete Banach lattices.
The second main goal of this paper, and the one that motivates the first, is to settle this representation problem in the negative. We exhibit unital Banach lattice algebras of dimension $2$ that admit no faithful representation—unital or not—on any Dedekind complete vector lattice (Example 4.4). This answers [undefh, Question 5.3] and [undefi, Problem 2] in the negative.
Wickstead studied in [undefj] the representability of all unital two-dimensional Banach lattice algebras. He noted that, in $\mathbb{R}^{2}$ with the pointwise product, there exists a family of lattice cones for which the resulting Banach lattice algebras admit a unital representation if and only if $\alpha$ takes all values between 0 and $1/2$. When $X$ is finite-dimensional, Wickstead showed that the only values $\alpha$ can possibly take are $\alpha=0$ or $\alpha=1/n$, where $n$ is a divisor of the dimension of $X$. It was left open whether $\alpha$ can take any other values or not. In 3.1 we show that it cannot. The unital two-dimensional examples announced above are then the cones in Wickstead’s family corresponding to the forbidden values of $\alpha$.
The result on positive projections will follow from a version of it for positive projections with constant diagonal on general $C(K)$ spaces (2.1). As a consequence, we will derive a condition guaranteeing that a Banach lattice algebra does not admit any faithful representation—unital or not—on a space of regular operators on a Dedekind complete Banach lattice (4.3). We will also generalize Wickstead’s complete description of contractive positive projections with constant diagonal on finite-dimensional Banach lattices to $\ell_{p}(A)$, for an arbitrary set $A$ and $1\leq p<\infty$ (3.2).
1.1. Preliminaries and notation
For a compact Hausdorff space $K$, we denote the space of continuous functions on $K$ by $C(K)$ and its dual, which we identify with the space of all finite signed regular Borel measures on $K$, by $M(K)$. For every $f\in C(K)$ and $\mu\in M(K)$, we denote the dual pairing by $\left<\mu,f\right$>, whilst the value of $\mu$ at a Borel set $B$ will be denoted by $\mu(B)$. The point mass measure at $t\in K$ will be denoted by $\delta_{t}$. Every $f\in C(K)$ has an associated multiplication operator $M_{f}\colon C(K)\to C(K)$ defined by $M_{f}(g)=fg$, for all $g\in C(K)$.
For the basic theory of vector and Banach lattices, we refer the reader to the monographs [undefg, undeff]. For the theory of positive operators and related topics, see [undef, undefa]; in particular, a detailed treatment of positive projections can be found in [undef, Chapter 5].
Now we want to introduce a couple of definitions that are convenient for this paper and, in order to do so, we have to recall first some basic facts of positive and regular operators. Let $X$ be a vector lattice. An operator $T\colon X\to X$ is central if there exists $\lambda>0$ such that $\pm T\leq\lambda\operatorname{id}_{X}$, where $\operatorname{id}_{X}$ denotes the identity of $X$. The center of $X$ is the set formed by all central operators, and it is denoted by $\mathcal{Z}(X)$. The central operators on $C(K)$ are precisely the multiplication operators. When $X$ is Dedekind complete, the space of regular operators $\mathcal{L}_{r}(X)$ becomes a Dedekind complete vector lattice under the Riesz–Kantorovich formulas. In this case, $\mathcal{Z}(X)$ is the ideal generated by $\operatorname{id}_{X}$, and it is in fact a band projection. The projection onto this band $\mathcal{D}\colon\mathcal{L}_{r}(X)\to\mathcal{L}_{r}(X)$ is usually called the diagonal projection or the projection onto the diagonal. For this reason, we introduce the following terminology.
Definition 1.1.
Let $X$ be a Dedekind complete vector lattice, and let $T\colon X\to X$ be a regular operator. We say that $T$ has constant diagonal $\alpha$ if $\mathcal{D}(T)=\alpha\operatorname{id}_{X}$.
Notice that $T$ having constant diagonal $\alpha$ is the same as $\alpha\operatorname{id}_{X}$ being the supremum of all the central operators that lie below $T$. This is the characterization we used at the beginning of the introduction.
A lattice-ordered algebra is a vector lattice together with a real algebra structure for which the product of positive elements is positive. The monograph [undefb, Chapter XVII] and the surveys [undefc, undefe] provide an overview of lattice-ordered rings and algebras. A Banach lattice algebra is a lattice-ordered algebra together with a complete lattice norm that is submultiplicative. An introduction to the theory of Banach lattice algebras can be found in [undefh].
Definition 1.2.
A representation of a lattice-ordered algebra $A$ is a lattice and algebra homomorphism $\Phi\colon A\to\mathcal{L}_{r}(X)$, for some Dedekind complete vector lattice $X$. We say that $\Phi$ is faithful if it is injective. In the case that $A$ has an algebraic identity $e$ and $\Phi(e)=\operatorname{id}_{X}$, we say that $\Phi$ is unital.
2. The main theorem
This section is devoted to proving the following.
Theorem 2.1.
Let $K$ be a non-empty compact Hausdorff space. Let $P\colon C(K)\to C(K)$ be a positive projection with $P\mathbb{1}_{K}=\mathbb{1}_{K}$. Suppose there exists some $\alpha\geq 0$ satisfying:
  1. (i)
    $\alpha\operatorname{id}_{C(K)}\leq P$, and
  2. (ii)
    if $0\leq M_{f}\leq P$, then $\|f\|_{\infty}\leq\alpha$.
Then either $\alpha=0$ or $\alpha=1/n$ for some $n\in\mathbb{N}$.
We will construct the proof incrementally along this section. We start with some lemmas that may be of interest by themselves.
Lemma 2.2.
Let $X$ be a Banach lattice, and let $P\colon X\to X$ be a positive projection. Let $S\subseteq X_{+}$ be a set, and let $I$ be the ideal generated by $S$. If $P(S)\subseteq I$, then also $P(I)\subseteq I$ and $P(\bar{I})\subseteq\bar{I}$.
Proof.
If $x\in I$, then there exist $s_{1},\ldots,s_{n}\in S$ and $\lambda_{1},\ldots,\lambda_{n}\in\mathbb{R}_{+}$ such that
$|x|\leq\lambda_{1}s_{1}+\cdots+\lambda_{n}s_{n}.$
Using the positivity of $P$ it follows that
$|Px|\leq P|x|\leq\lambda_{1}Ps_{1}+\cdots+\lambda_{n}Ps_{n}.$
Since $P(S)\subseteq I$, the right hand side in previous equation belongs to $I$, and therefore $Px\in I$. Thus $P(I)\subseteq I$ and, since $P$ is continuous, $P(\bar{I})\subseteq\bar{I}$. ∎
Lemma 2.3.
Let $X$ be a Banach lattice with a quasi-interior point. Let $P\colon X\to X$ be a positive projection satisfying $\alpha\operatorname{id}_{X}\leq P$, for some $\alpha>0$. If $B\subseteq X$ is a projection band invariant under $P$, then so is $B^{d}$.
Proof.
Let $e$ be a quasi-interior point of $X$, and let $f=Pe$. Since $f=Pe\geq\alpha e$, and $\alpha>0$, $f$ is also a quasi-interior point. Let $Q\colon X\to X$ be the band projection onto $B$. Then using the positivity of $P$, and the fact that $Qf\leq f$, it follows that
$PQf\leq Pf=f.$
Since $PQf\in B$, applying $Q$ to this equality yields $PQf\leq Qf$. Thus
$0=P(Qf-PQf)\geq\alpha(Qf-PQf)\geq 0$
and the equality $PQf=Qf$ follows. Using this, the computation
$(\operatorname{id}_{X}-Q)f=f-Qf=Pf-PQf=P(\operatorname{id}_{X}-Q)f$
shows that $(\operatorname{id}_{X}-Q)f$ belongs to the image of $P$. By 2.2, the closed ideal generated by $(\operatorname{id}_{X}-Q)f$ is invariant under $P$. Since $f$ is a quasi-interior point, this closed ideal is precisely $B^{d}$. Indeed, for every $x\in X$, find a sequence $(x_{n})$ contained in the ideal generated by $f$ and such that $x_{n}\to x$ (this we can do because $f$ is a quasi-interior point). Then $(\operatorname{id}_{X}-Q)x_{n}\to(\operatorname{id}_{X}-Q)x$, where the sequence $((\operatorname{id}_{X}-Q)x_{n})$ belongs to the ideal generated by $(\operatorname{id}_{X}-Q)f$. ∎
Recall that a projection $P\colon X\to X$ on a vector lattice $X$ is said to be strictly positive if $x>0$ implies $Px>0$.
Lemma 2.4.
Let $X$ be a Banach lattice, and let $P\colon X\to X$ be a strictly positive projection. If the only closed ideals invariant under $P$ are the trivial ones (that is, $\{0\}$ and $X$), then $P$ has rank $1$.
Proof.
Since $P$ is strictly positive, by [undef, Lemma 5.43] $P(X)$ is a (closed) vector sublattice of $X$. If $P(X)$ didn’t have dimension $1$, then there would exist two positive disjoint vectors $x,y\in P(X)$. The closed ideal generated by $x$ in $X$ is non-trivial (because $y$ does not belong to the band generated by $x$), and yet invariant by 2.2. This contradicts the assumption. ∎
Let $K$ be a compact Hausdorff space, and let $P\colon C(K)\to C(K)$ and $\alpha\geq 0$ be as in 2.1. If $\alpha=0$ we are done, so we may assume from now on that $\alpha>0$. The adjoint $P^{*}\colon M(K)\to M(K)$ is again a positive projection satisfying $\alpha\operatorname{id}_{M(K)}\leq P^{*}$. For every $t\in K$, let $\mu_{t}=P^{*}\delta_{t}$. Notice that the map
[…]
is continuous for the weak topology on $M(K)$; this fact will be used several times later on.
Decompose
$\mu_{t}=\sum_{s\in J_{t}}\lambda(t,s)\delta_{s}+\nu_{t},$
where $\nu_{t}$ is a positive continuous measure (i.e., it is disjoint from every point mass measure), $J_{t}\subseteq K$ is an at most countable set of points, and $\lambda(t,s)>0$ for every $s\in J_{t}$. Moreover, since $\alpha\operatorname{id}_{M(K)}\leq P^{*}$ and $\alpha>0$, we must have $t\in J_{t}$ and $\alpha\leq\lambda(t,t)$ for all $t\in K$. Also, it is important to keep in mind that $\mu_{s}$ is positive and
$\|\mu_{s}\|=\mu_{s}(K)=\left<P^{*}\delta_{s},\mathbb{1}_{K}\right>=\left<\delta_{s},P\mathbb{1}_{K}\right>=\left<\delta_{s},\mathbb{1}_{K}\right>=1.$
Lemma 2.5.
For every $s\in J_{t}$, $J_{s}\subseteq J_{t}$.
Proof.
Using that $P^{*}$ is a projection, we have
(1) $\sum_{s\in J_{t}}\lambda(t,s)\delta_{s}+\nu_{t}=\mu_{t}=P^{*}\mu_{t}=\sum_{s\in J_{t}}\lambda(t,s)\mu_{s}+P^{*}\nu_{t}.$
Notice that all the coefficients are positive. Hence, for this equality to hold, the discrete part of each $\mu_{s}$ cannot contain any point mass measure supported outside $J_{t}$. ∎
Fix an arbitrary $t\in K$. Denote by $I$ the order ideal generated by $\{\,\mu_{s}:s\in J_{t}\,\}$ in $M(K)$. Since $P^{*}\mu_{s}=\mu_{s}$ for every $s\in J_{t}$, 2.2 guarantees that both $I$ and $\bar{I}$ are invariant under $P^{*}$. Since $M(K)$ is order continuous, $\bar{I}$ is actually a projection band.
Lemma 2.6.
The only closed invariant ideals of $P^{*}|_{\bar{I}}$ are the trivial ones.
Proof.
We want to use 2.3. For this, first note that $\bar{I}$, being a band in $M(K)$, is itself an order continuous Banach lattice (actually, it is an AL-space). Also note that $\bar{I}$ has a quasi-interior point: if $J_{t}=\{t_{i}\}_{i=1}^{\infty}$ is an enumeration of $J_{t}$, then $e=\sum_{i=1}^{\infty}2^{-i}\mu_{t_{i}}$ is a quasi-interior point of $\bar{I}$ (recall that $\|\mu_{s}\|=1$).
Suppose $I^{\prime}\subseteq\bar{I}$ is a closed ideal invariant under $P^{*}|_{\bar{I}}$. The order continuity of $\bar{I}$ implies that $I^{\prime}$ is a projection band in $I^{\prime}$. By 2.3, its disjoint complement in $\bar{I}$ is also invariant under $P^{*}|_{\bar{I}}$. Now $\delta_{t}\in\bar{I}$ is an atom, and therefore must belong to either $I^{\prime}$ or its disjoint complement. Without loss of generality, assume it belongs to $I^{\prime}$. Then the invariance of $I^{\prime}$ under $P^{*}$ implies that
$P^{*}\delta_{t}=\mu_{t}=\sum_{s\in J_{t}}\lambda(t,s)\delta_{s}+\nu_{t}\in I^{\prime}.$
Since $\lambda(t,s)>0$ for every $s\in J_{t}$, and $I^{\prime}$ is an ideal, $\delta_{s}\in I^{\prime}$ and, using again invariance, $\mu_{s}\in I^{\prime}$ for all $s\in J_{t}$. But $\bar{I}$ is precisely the closed ideal generated by $\mu_{s}$, so that $I^{\prime}=\bar{I}$. ∎
By 2.4, $P^{*}(\bar{I})$ has rank one. That is, there exist a measure $\rho\in M(K)_{+}$ and scalars $\lambda_{s}\in\mathbb{R}_{+}$ such that $\mu_{s}=P^{*}\mu_{s}=\lambda_{s}\rho$ for all $s\in J_{t}$. But since $\mu_{s}(K)=1$ for every $s\in J_{t}$, $\lambda_{s}$ does not depend on $s$, and therefore $\mu_{s}=\mu_{t}$ for every $s\in J_{t}$. This implies that the family of all distinct sets $J_{t}$ forms a partition of $K$.
Lemma 2.7.
For all $s\in K$, either $J_{s}=J_{t}$ (when $s\in J_{t}$) or $J_{s}\cap J_{t}=\varnothing$ (when $s\not\in J_{t}$).
Proof.
When $s\in J_{t}$, we have noted already that $\mu_{s}=\mu_{t}$, so obviously $J_{s}=J_{t}$. If $s^{\prime}\in J_{s}\cap J_{t}$, then again we would have $\mu_{s}=\mu_{s^{\prime}}=\mu_{t}$, and therefore $s\in J_{s}=J_{t}$. ∎
The fact that $\mu_{s}=\mu_{t}$ for all $s\in J_{t}$ has further implications on the expression of $\mu_{t}$. First, since the continuous parts of $\mu_{t}$ and $\mu_{s}$ must be equal, $\nu_{t}=\nu_{s}$ holds for every $s\in J_{t}$. Second, since the coefficient of each point evaluation must be equal, $\lambda(t,s)=\lambda(s,s)$ holds for every $s\in J_{t}$. Hence $\lambda(t,s)\geq\alpha$ for every $s\in J_{t}$. Moreover, since $\alpha>0$ and $\|\mu_{s}\|=1$, $J_{t}$ must be finite and satisfy $|J_{t}|\leq 1/\alpha$.
Next we want to show that $\nu_{t}=0$. For this we need the following topological lemma.
Lemma 2.8.
For every $\varepsilon>0$ and $t^{\prime}\in J_{t}$, there exists an open neighborhood $V$ of $t^{\prime}$ such that $(P^{*}\nu_{t})(V)<\varepsilon$.
Proof.
Let $J_{t}=\{t_{1},\ldots,t_{n}\}$, and let $U_{1}^{\prime},\ldots,U_{n}^{\prime}$ be pairwise disjoint open neighborhoods of $t_{1},\ldots,t_{n}$, respectively. Let $U^{\prime}=U_{1}^{\prime}\cup\cdots\cup U_{n}^{\prime}$. Since $\nu_{t}$ is a regular continuous measure, we may choose these neighborhoods small enough so as to have $\nu_{t}(U^{\prime})<\varepsilon/2$. Let also $U_{i}$ be an open neighborhood of $t_{i}$ satisfying $\overline{U_{i}}\subseteq U_{i}^{\prime}$ for $i=1,\ldots,n$, and denote $U=U_{1}\cup\cdots\cup U_{n}$.
Fix $t^{\prime}\in J_{t}$. For every $s\not\in U$, $J_{s}$ does not contain any of the points in $J_{t}$ by 2.7. Therefore there exists an open neighborhood $V_{s}$ of $t^{\prime}$ such that $\mu_{s}(V_{s})<\varepsilon/2$. Recall that the map $s^{\prime}\mapsto\mu_{s^{\prime}}$ is continuous in the weak topology. Hence there exists an open neighborhood $W_{s}$ of $s$ such that $\mu_{s^{\prime}}(V_{s})<\varepsilon/2$ for all $s^{\prime}\in W_{s}$.
This way the $\{W_{s}\}_{s\in K\setminus U}$ form an open covering of the set $K\setminus U$. Since the latter is compact, we can extract a finite subcovering:
$K\setminus U\subseteq W_{s_{1}}\cup\cdots\cup W_{s_{m}}.$
Consider the set $V=V_{s_{1}}\cap\cdots\cap V_{s_{m}}$, which is again an open neighborhood of $t^{\prime}$. Notice that, if $s\not\in U$, then $s\in W_{s_{i}}$ for some $i\in\{1,\ldots,m\}$, and therefore
$\mu_{s}(V)\leq\mu_{s}(V_{s_{i}})<\varepsilon/2.$
To prove that $V$ is the set we wanted to construct, we need to show that, for every $f\in C(K)$ satisfying $0\leq f\leq\mathbb{1}_{K}$ and $\operatorname{supp}f\subseteq V$, we have
$\left<P^{*}\nu_{t},f\right>=\left<\nu_{t},Pf\right><\varepsilon.$
Fix such an $f\in C(K)$. Let $U_{1}=U^{\prime}$, $U_{2}=K\setminus\bar{U}$, and let $\{f_{1},f_{2}\}$ be a partition of unit subordinate to the open covering $\{U_{1},U_{2}\}$. On the one hand, for every $s\not\in U$:
$(Pf)(s)f_{2}(s)\leq\left<P^{*}\delta_{s},f\right>=\left<\mu_{s},f\right>\leq\mu_{s}(V)<\varepsilon/2.$
Since $f_{2}(s)=0$ for $s\in U$, this shows that $(Pf)f_{2}<\varepsilon/2\mathbb{1}_{K}$. In particular,
$\left<\nu_{t},(Pf)f_{2}\right>\leq\varepsilon/2\nu_{t}(K)\leq\varepsilon/2.$
On the other hand, since $\|Pf\|_{\infty}\leq\|f\|_{\infty}\|P\mathbb{1}_{K}\|_{\infty}\leq 1$,
$\left<\nu_{t},(Pf)f_{1}\right>\leq\left<\nu_{t},f_{1}\right>\leq\nu_{t}(U^{\prime})<\varepsilon/2.$
Therefore
$\left<\nu_{t},Pf\right>=\left<\nu_{t},(Pf)f_{1}\right>+\left<\nu_{t},(Pf)f_{2}\right><\varepsilon.\qed$
Lemma 2.9.
For every $t\in K$, $\nu_{t}=0$.
Proof.
Expanding the right hand side of (1), we have the following identity:
(2) $\sum_{s\in J_{t}}\lambda(t,s)\delta_{s}+\nu_{t}=\sum_{s\in J_{t}}\sum_{s^{\prime}\in J_{t}}\lambda(t,s)\lambda(s,s^{\prime})\delta_{s^{\prime}}+\sum_{s\in J_{t}}\lambda(t,s)\nu_{t}+P^{*}\nu_{t},$
where we are using that $J_{s}=J_{t}$ and also that $\nu_{s}=\nu_{t}$ for every $s\in J_{t}$. Fix $t^{\prime}\in J_{t}$. According to 2.8, for every $\varepsilon>0$ there exists an open neighborhood $V$ of $t^{\prime}$ such that $(P^{*}\nu_{t})(V)<\varepsilon$. By making $V$ smaller, we can assume that also $\nu_{t}(V)<\varepsilon$ and that $t^{\prime\prime}\not\in V$ for every other $t^{\prime\prime}\in J_{t}$ different from $t^{\prime}$. Evaluating the above expression at $V$ yields
$\lambda(t,t^{\prime})+\nu_{t}(V)=\sum_{s\in J_{t}}\lambda(t,s)\lambda(s,t^{\prime})+\sum_{s\in J_{t}}\lambda(t,s)\nu_{t}(V)+(P^{*}\nu_{t})(V).$
Using that $\sum_{s\in J_{t}}\lambda(t,s)\leq 1$, rearrange the previous expression to get
$\bigg|\lambda(t,t^{\prime})-\sum_{s\in J_{t}}\lambda(t,s)\lambda(s,t^{\prime})\bigg|<3\varepsilon.$
Since $\varepsilon>0$ was arbitrary, it follows that $\lambda(t,t^{\prime})=\sum_{s\in J_{t}}\lambda(t,s)\lambda(s,t^{\prime})$.
Using these identities for every $t^{\prime}\in J_{t}$, (2) becomes
$\nu_{t}=\sum_{s\in J_{t}}\lambda(t,s)\nu_{t}+P^{*}\nu_{t}.$
Applying $P^{*}$ to the above equation, and the fact that $\sum_{s\in J_{t}}\lambda(t,s)>0$, yields $P^{*}\nu_{t}=0$. Therefore
$\nu_{t}=\sum_{s\in J_{t}}\lambda(t,s)\nu_{t}$
and either $\nu_{t}=0$, in which case we are done, or $\sum_{s\in J_{t}}\lambda(t,s)=1$. But since
$1=\mu_{t}(K)=\sum_{s\in J_{t}}\lambda(t,s)+\nu_{t}(K),$
in the latter case we also end up having $\nu_{t}=0$. ∎
So far we have proved that, for every $t\in K$,
$\mu_{t}=\sum_{s\in J_{t}}\lambda(t,s)\delta_{s},$
where $\alpha\leq\lambda(t,s)$ and $J_{t}$ is finite with uniformly bounded size $|J_{t}|\leq 1/\alpha$. Up to this point we have not used that $\alpha\operatorname{id}_{C(K)}$ is the biggest multiplication operator below $P$, only that $\alpha\operatorname{id}_{C(K)}\leq P$. Now this fact has to come into play.
Lemma 2.10.
For every $n\in\mathbb{N}$, there exists a dense set $D_{n}\subseteq K$ such that $\lambda(t,t)\leq\alpha+1/n$ for all $t\in D_{n}$.
Proof.
Let $n\in\mathbb{N}$. In order to get a contradiction, suppose there exists an open set $U\subseteq K$ such that $\lambda(t,t)>\alpha+1/n$ for all $t\in U$. Let $V$ be an open set with $\bar{V}\subseteq U$. By Urysohn’s lemma, there exists a continuous function $f\in C(K)$ such that $0\leq f\leq\mathbb{1}_{K}$, $\operatorname{supp}f\subseteq U$, and $f|_{V}=1$.
Set $g=(\alpha+1/n)f$. Then for every $h\in C(K)_{+}$ and $t\in U$:
$g(t)h(t)\leq(\alpha+1/n)h(t)\leq\lambda(t,t)h(t)\leq(Ph)(t).$
Also, $g(t)h(t)=0\leq(Ph)(t)$ whenever $t\not\in U$. Hence $M_{g}\leq P$, but $\|g\|_{\infty}=\alpha+1/n>\alpha$, contradicting the assumption. ∎
We can now use this fact to obtain the final lemma.
Lemma 2.11.
There exists $t\in K$ such that $\lambda(t,s)=\alpha$ for all $s\in J_{t}$.
Proof.
Since $|J_{s}|\leq 1/\alpha$ for all $s\in K$, we can choose $t\in K$ for which the cardinality of $J_{t}$ is maximal. We claim that $\lambda(t,s)=\alpha$ for all $s\in J_{t}$. To show this, we will proceed by contradiction: suppose there exist some $t_{1}\in J_{t}$ and some $n\in\mathbb{N}$ for which $\lambda(t,t_{1})>\alpha+1/n$.
Since $J_{t_{1}}=J_{t}$, $J_{t_{1}}$ also has maximal cardinality. Write $J_{t}=\{t_{1},\ldots,t_{m}\}$, and let $U_{1},\ldots,U_{m}$ be pairwise disjoint open neighborhoods of $t_{1},\ldots,t_{m}$. By the weak continuity of $s\mapsto\mu_{s}$, there exists an open neighborhood $U$ of $t_{1}$ such that, whenever $s\in U$, $\mu_{s}(U_{1})>\alpha+1/n$ (recall that $\mu_{t_{1}}(U_{1})>\alpha+1/n$) and $\mu_{s}(U_{i})>\alpha/2$ (recall that $\mu_{t_{1}}(U_{i})\geq\alpha$) for $i=2,\ldots,m$. This means that, for every $s\in U$, $J_{s}$ must contain at least a point in each $U_{1},\ldots,U_{m}$. Since these sets are pairwise disjoint, $J_{s}$ contains at least $m$ points. The maximality of $m$ implies then that $J_{s}$ must contain exactly $m$ points, one in each of the sets $U_{1},\ldots,U_{m}$.
In particular, for the $s\in U\cap U_{1}$, the point of $J_{s}$ that belongs to $U_{1}$ is precisely $s$. And since $\mu_{s}(U_{1})>\alpha+1/n$, and the only point of $J_{s}$ that belongs to $U_{1}$ is $s$, it follows that $\lambda(s,s)>\alpha+1/n$ for every $s\in U\cap U_{1}$. This contradicts 2.10. ∎
Pick $t\in K$ as in the lemma above. Then
$\mu_{t}=\sum_{s\in J_{t}}\alpha\delta_{s}$
so that
$1=\mu_{t}(K)=\sum_{s\in J_{t}}\alpha=\alpha|J_{t}|$
which implies $\alpha=1/|J_{t}|$. This completes the proof of 2.1.
We conclude this section with a general example showing how each of the values 0 and $1/n$, for $n\in\mathbb{N}$, can be attained.
Example 2.12.
Let $K$ be a compact Hausdorff space. Let $G$ be a finite subgroup of the homeomorphism group of $K$. Then the map $P\colon C(K)\to C(K)$ defined by
$Pf=\frac{1}{|G|}\sum_{\sigma\in G}f\circ\sigma,\quad\text{for }f\in C(K),$
is a positive projection that satisfies the conditions of 2.1 with $\alpha=1/|G|$. Indeed, it is immediate to check that $P$ is a positive projection such that $1/|G|\operatorname{id}_{C(K)}\leq P$ and $P\mathbb{1}_{K}=\mathbb{1}_{K}$. Suppose $0\leq M_{h}\leq P$ for some $h\in C(K)_{+}$. Given $t\in K$, choose an open neighborhood $U$ of $t$ such that $\sigma(t)\not\in U$ for all $\sigma\in G\setminus\{\operatorname{id}_{K}\}$, and let $f\in C(K)$ be a continuous function satisfying $f(t)=1$, $0\leq f\leq\mathbb{1}_{K}$, and $\operatorname{supp}f\subseteq U$. Then
$h(t)=M_{h}(f)(t)\leq(Pf)(t)=\frac{1}{|G|}.$
Since this is true for every $t\in K$, $\|h\|_{\infty}\leq 1/|G|$.
To construct an example with $\alpha=0$, let $\mu\in M(K)_{+}$ be a continuous measure with $\mu(K)=1$, and define
$Pf=\mathbb{1}_{K}\int_{K}f\,d\mu\quad\text{for }f\in C(K).$
It is immediate to check that $P$ is a positive projection satisfying $P\mathbb{1}_{K}=\mathbb{1}_{K}$. Suppose $0\leq M_{h}\leq P$, for some $h\in C(K)_{+}$. Fix $t\in K$ and $\varepsilon>0$. Let $U$ be an open neighborhood of $t$ for which $\mu(U)\leq\varepsilon$. Let $f\in C(K)$ be a continuous function satisfying $f(t)=1$, $0\leq f\leq\mathbb{1}_{K}$, and $\operatorname{supp}f\subseteq U$. Then
$h(t)=M_{h}(f)(t)\leq(Pf)(t)\leq\mu(U)<\varepsilon.$
Since $\varepsilon>0$ and $t\in K$ were arbitrary, $h=0$.
3. Wickstead’s conjecture and other consequences
2.1 can now be used to answer [undefj, Question 7.3].
Corollary 3.1.
Let $X$ be a non-zero Dedekind complete vector lattice. Let $P\colon X\to X$ be a positive projection with constant diagonal $\alpha$. Then either $\alpha=0$ or $\alpha=1/n$, for some $n\in\mathbb{N}$.
Proof.
If $\alpha=0$ we are done, so assume $\alpha>0$. In particular, $P\neq 0$, so there exists some $e\in X_{+}$ such that $Pe\neq 0$. Then $I_{Pe}$, the ideal generated by $Pe$, is invariant under $P$ by 2.2. Use Kakutani’s theorem [undefd, Theorem 1] to identify $(I_{Pe},\|{\cdot}\|_{Pe})$ lattice isometrically with $C(K)$, for some non-empty compact Hausdorff space $K$, in such a way that $Pe$ is identified with the function $\mathbb{1}_{K}$. We are going to show that the restriction of $P$ to $I_{Pe}$ satisfies the conditions of 2.1.
It is clear that the restriction is a positive projection satisfying $\alpha\operatorname{id}_{C(K)}\leq P$. Moreover, $P\mathbb{1}_{K}=P(Pe)=Pe=\mathbb{1}_{K}$. Since $X$ is Dedekind complete, and the diagonal of $P$ is $\alpha\operatorname{id}_{X}$, we can express $P=\alpha\operatorname{id}_{X}+T$, where $T$ is a positive operator disjoint from $\operatorname{id}_{X}$. By the Riesz–Kantorovich formulas [undefa, Theorem 1.18], this means that
$0=(T\wedge\operatorname{id}_{X})(x)=\inf_{0\leq y\leq x}Ty+(x-y)$
holds for every $x\in X_{+}$. Note that $T=P-\alpha\operatorname{id}_{X}$ also leaves $I_{Pe}$ invariant. Moreover, since $I_{Pe}$ is itself a Dedekind complete Banach lattice, the Riesz–Kantorovich formulas also hold in $\mathcal{L}_{r}(I_{Pe})$. Therefore, for every $x\in(I_{Pe})_{+}$,
$(T|_{I_{Pe}}\wedge\operatorname{id}_{X})(x)=\inf_{y\in I_{Pe},0\leq y\leq x}Ty+(x-y)=\inf_{0\leq y\leq x}Ty+(x-y)=0,$
where in the second equality we are using that $I_{Pe}$ is an ideal. Hence, in $\mathcal{L}_{r}(I_{Pe})$, $P|_{I_{Pe}}=\alpha\operatorname{id}_{I_{Pe}}+T|_{I_{Pe}}$, with $T$ disjoint from the identity. Now, if $f\in C(K)_{+}$ is such that $M_{f}\leq P$, then because of the previous decomposition and the fact that $M_{f}$ is a central operator $M_{f}\leq\alpha\operatorname{id}_{C(K)}$. Hence $\|f\|_{\infty}\leq\alpha$. ∎
Previous result applies, in particular, to conditional expectations with constant diagonal. We can use this fact, together with the ideas in the proof of 2.1, to characterize all contractive positive projections with constant diagonal on $\ell_{p}$ spaces. Given a set $A$, $\chi_{B}$ will denote the characteristic function of $B\subseteq A$, and $e_{a}$ will denote the characteristic function of the singleton $\{a\}\subseteq A$.
Proposition 3.2.
Let $A$ be a non-empty set, let $1\leq p<\infty$, and let $P\colon\ell_{p}(A)\to\ell_{p}(A)$ be a contractive positive projection with constant diagonal $\alpha>0$. Then $n=1/\alpha\in\mathbb{N}$, and there exists a partition $\{A_{i}\}_{i}$ of $A$ into sets of size $n$ such that
$P\bigg(\sum_{a\in A}\lambda_{a}e_{a}\bigg)=\alpha\sum_{i}\bigg(\sum_{a\in A_{i}}\lambda_{a}\bigg)\chi_{A_{i}},$
whenever $\sum_{a\in A}|\lambda_{a}|^{p}<\infty$. In particular, if $A$ is countable, then there exists an enumeration of $A$ such that the matrix associated with $P$ is
$\begin{pmatrix}\alpha\bm{1}_{n}&\bm{0}_{n}&\bm{0}_{n}&\cdots\\ \bm{0}_{n}&\alpha\bm{1}_{n}&\bm{0}_{n}&\cdots\\ \bm{0}_{n}&\bm{0}_{n}&\alpha\bm{1}_{n}&\cdots\\ \vdots&\vdots&\vdots&\ddots\end{pmatrix}$,
where $\bm{0}_{n}$ denotes an $n$ by $n$ block of zeros, and $\bm{1}_{n}$ denotes an $n$ by $n$ block of ones.
Proof.
For every $a\in A$, we can express
$Pe_{a}=\sum_{b\in J_{a}}\lambda(a,b)e_{b},$
where $\lambda(a,b)>0$ and $J_{a}\subseteq A$ is at most countable. Moreover, that the diagonal of $P$ is $\alpha$ implies $(Pe_{a})(a)=\alpha$, that is, $a\in J_{a}$ and $\lambda(a,a)=\alpha$.
Fix $a\in A$. We claim that, for every $b\in J_{a}$, $J_{b}\subseteq J_{a}$. Indeed, since $P$ is a projection,
$\sum_{b\in J_{a}}\lambda(a,b)e_{b}=Pe_{a}=P(Pe_{a})=\sum_{b\in J_{a}}\lambda(a,b)\sum_{c\in J_{b}}\lambda(b,c)e_{c}.$
Since all the coefficients are positive, it cannot be the case that $J_{b}$, for $b\in J_{a}$, introduces a new coordinate. Hence $J_{b}\subseteq J_{a}$.
Let $I$ be the closed linear span of $\{\,e_{b}:b\in J_{a}\,\}$. Notice that $I$ is a closed ideal, and therefore a band. By the above observation and 2.2, this ideal is invariant under $P$. We are going to show that the restriction $P|_{I}\colon I\to I$ has rank $1$ just like in 2.6. By 2.4, it suffices to show that the only invariant closed ideals are the trivial ones. Let $I^{\prime}\subseteq I$ be a closed ideal invariant under $P|_{I}$. Since $I$ is order continuous and has a quasi-interior point (in fact, it is lattice isometric to $\ell_{p}$), it follows from 2.3 that the disjoint complement of $I^{\prime}$ in $I$ is also invariant. But $e_{a}$ is an atom: it must belong to either $I^{\prime}$ or its disjoint complement. Suppose, without loss of generality, that $e_{a}\in I^{\prime}$. Then by invariance $Pe_{a}\in I^{\prime}$ and, being $I^{\prime}$ an ideal, $e_{b}\in J_{a}$ for all $b\in J_{a}$. Since $I$ is the smallest closed ideal containing all the $e_{b}\in J_{a}$, it follows that $I^{\prime}=I$. Hence $P|_{I}$ has rank one.
Let $x$ be a positive norm-one element in $P(I)$. Notice that, in particular, $x(b)>0$ for all $b\in J_{a}$. Define a probability measure $\mu$ on $J_{a}$ by $\mu(b)=x(b)^{p}$, for $b\in J_{a}$. Then $I$ is lattice isometric to $L_{p}(\mu)$ through the map
$\begin{array}[]{cccc}J\colon&I&\longrightarrow&L_{p}(\mu)\\ &y&\longmapsto&(y(b)x(b)^{-1})_{b\in J_{a}}\\ \end{array}.$
Define $\tilde{P}=JP|_{I}J^{-1}\colon L_{p}(\mu)\to L_{p}(\mu)$. This is still a contractive positive projection, and it satisfies
$\tilde{P}(\mathbb{1}_{J_{a}})=JPJ^{-1}(\mathbb{1}_{J_{a}})=JPx=Jx=\mathbb{1}_{J_{a}}.$
By the Andô–Douglas theorem [undef, Corollary 5.53], $\tilde{P}$ is the conditional expectation with respect to some $\sigma$-algebra $\Sigma$ on $J_{a}$. Since $J_{a}$ is at most countable, $\Sigma$ must be generated by a partition $\{A_{i}\}$ of $J_{a}$. Hence the conditional expectation has the form
$\tilde{P}(f)=\sum_{i}\chi_{A_{i}}\frac{1}{\mu(A_{i})}\sum_{b\in A_{i}}\mu(b)f(b)\quad\text{for all }f\in L_{p}(\mu).$
From $\tilde{P}(L_{p}(\mu))=\operatorname{span}\{\mathbb{1}_{J_{a}}\}$ it follows that the partition must be trivial, so that actually
$\tilde{P}(f)=\mathbb{1}_{J_{a}}\sum_{b\in J_{a}}\mu(b)f(b)\quad\text{for all }f\in L_{p}(\mu),$
and therefore
$P|_{I}(y)=J^{-1}\tilde{P}Jy=J^{-1}(\mathbb{1}_{J_{a}})\sum_{b\in J_{a}}\mu(b)x(b)^{-1}y(b)=x\sum_{b\in J_{a}}x(b)^{p-1}y(b)\quad\text{for all }y\in I.$
In particular, $\alpha=(Pe_{b})(b)=x(b)^{p}$ for all $b\in J_{a}$. Since $\alpha>0$, and $\|x\|_{p}=1$, this means that $J_{a}$ must be finite with $|J_{a}|=1/\alpha$. Also, the expression for $P|_{I}$ ends up being
(3) $P|_{I}(y)=\chi_{J_{a}}\alpha\sum_{b\in J_{a}}y(b)\quad\text{for all }y\in I.$
Therefore, for every $b\in J_{a}$, $Pe_{b}=\alpha\chi_{J_{a}}=Pe_{a}$. Hence $J_{a}=J_{b}$ whenever $b\in J_{a}$, and $J_{a}\cap J_{b}=\varnothing$ otherwise.
Consider then the partition $\{A_{i}\}$ formed by the sets of the form $J_{a}$, $a\in A$, that are different. Using (3) we arrive at
$P\bigg(\sum_{a\in A}\lambda_{a}a\bigg)=P\bigg(\sum_{i}\sum_{a\in A_{i}}\lambda_{a}a\bigg)=\alpha\sum_{i}\bigg(\sum_{a\in A_{i}}\lambda_{a}\bigg)\chi_{A_{i}}$
whenever $\sum_{a\in A}|\lambda_{a}|^{p}<\infty$. In the case that $A$ is at most countable, to get the desired matrix first enumerate the partition $\{A_{k}\}_{k\in\mathbb{N}}$, and then enumerate $A$ in such a way that $A_{k}=\{n(k-1)+1,\ldots,nk\}$, where $n=1/\alpha$. ∎
Remark.
Following the ideas of previous proof, it is not difficult to check that the only contractive positive projection $P\colon\ell_{p}(A)\to\ell_{p}(A)$ with constant diagonal 0 is $P=0$.
4. On the representation of Banach lattice algebras
In this section we use 3.1 to settle the representation problem for Banach lattice algebras in the negative. We begin with two preparatory lemmas: the first records that the range of a positive projection is itself a Dedekind complete vector lattice, and the second shows that disjointness from a positive idempotent in $\mathcal{L}_{r}(X)$ becomes disjointness from the identity on its range. These lemmas will let us apply 3.1 on the range of an arbitrary positive idempotent of $\mathcal{L}_{r}(X)$.
In the following lemmas, when $X$ is a vector lattice, $Y$ is a sublattice, and $S\subseteq Y$ is a set, we denote by $\sup_{X}S$ the supremum of $S$ in $X$ (whenever it exists) and by $\sup_{Y}S$ the supremum of $S$ in $Y$ (whenever it exists).
Lemma 4.1.
Let $X$ be a Dedekind complete vector lattice, and let $E\colon X\to X$ be a positive projection. Then $Y=E(X)$, endowed with the order inherited from $X$, is a Dedekind complete vector lattice with supremum $\vee_{Y}$ given by
[…]
Proof.
Let $y_{1},y_{2}\in Y$. Since $E$ is positive and $Ey_{i}=y_{i}$ for $i=1,2$, we have $E(y_{1}\vee y_{2})\geq Ey_{i}=y_{i}$, so $E(y_{1}\vee y_{2})$ is an upper bound of $\{y_{1},y_{2}\}$ in $Y$. If $z\in Y$ satisfies $z\geq y_{1}$ and $z\geq y_{2}$, then $z\geq y_{1}\vee y_{2}$ in $X$, whence $z=Ez\geq E(y_{1}\vee_{X}y_{2})$. This shows that $y_{1}\vee_{Y}y_{2}=E(y_{1}\vee y_{2})$.
For Dedekind completeness, let $(y_{\gamma})$ be an increasing net in $Y$ that is bounded above by some $y_{0}\in Y$. By Dedekind completeness of $X$, $z=\sup_{X}y_{\gamma}$ exists. Then $Ez\geq Ey_{\gamma}=y_{\gamma}$ for all $\gamma$, so $Ez$ is an upper bound of $(y_{\gamma})$ in $Y$. If $w\in Y$ also satisfies $w\geq y_{\gamma}$ for all $\gamma$, then $w\geq z$ in $X$, and so $w=Ew\geq Ez$. Hence $Ez=\sup_{Y}y_{\gamma}$. ∎
Lemma 4.2.
Let $X$ be a Dedekind complete vector lattice, and let $E,T\in\mathcal{L}_{r}(X)$ be positive operators satisfying:
  1. (i)
    $E^{2}=E$,
  2. (ii)
    $ET=TE=T$, and
  3. (iii)
    $E\wedge T=0$ in $\mathcal{L}_{r}(X)$.
Set $Y=E(X)$, endowed with the order inherited from $X$. Then
$\operatorname{id}_{Y}\wedge T|_{Y}=0\quad\text{in }\mathcal{L}_{r}(Y).$
Proof.
By 4.1, $Y$ is a Dedekind complete vector lattice, so $\mathcal{L}_{r}(Y)$ is itself a Dedekind complete vector lattice and the Riesz–Kantorovich formulas hold in it. Note also that $T=ET=TE$ implies that $T$ leaves $Y$ invariant, so the restriction $T|_{Y}\colon Y\to Y$ is a well-defined element of $\mathcal{L}_{r}(Y)$.
Fix $y\in Y_{+}$. We compute $(E\wedge T)(y)$ using the Riesz–Kantorovich formula in $\mathcal{L}_{r}(X)$:
$(E\wedge T)(y)=\inf_{X}\{\,Ea+T(y-a):0\leq a\leq y\,\}.$
Since $T=TE$, for any $a\in X_{+}$ with $a\leq y$,
$Ea+T(y-a)=Ea+Ty-TEa=(\operatorname{id}_{Y}-T|_{Y})(Ea)+T|_{Y}(y).$
As $a$ ranges over $\{a\in X_{+}:a\leq y\}$, the element $w=Ea$ ranges over the order interval $0,y]_{Y}=\{w\in Y_{+}:w\leq y\}$ in $Y$. Therefore
(4) $(E\wedge T)(y)=\inf_{X}\{\,(\operatorname{id}_{Y}-T|_{Y})(w)+T|_{Y}(y):w\in[0,y]_{Y}\,\}.$
On the other hand, the Riesz–Kantorovich formula in $\mathcal{L}_{r}(Y)$ gives
(5) $(\operatorname{id}_{Y}\wedge T|_{Y})(y)=\inf_{Y}\big\{\,(\operatorname{id}_{Y}-T|_{Y})(w)+T|_{Y}(y):w\in[0,y]_{Y}\,\big\}.$
The set whose infimum is taken is the same subset of $Y$ in both (4) and (5). The only difference is that the first infimum is taken in $X$ while the second is taken in $Y\subseteq X$. Therefore
$(\operatorname{id}_{Y}\wedge T|_{Y})(y)\leq(E\wedge T)(y)=0.$
Since this holds for every $y\in Y_{+}$, we conclude that $\operatorname{id}_{Y}\wedge T|_{Y}=0$ in $\mathcal{L}_{r}(Y)$. ∎
We can now state the criterion for non-representability. The proposition below generalizes the criterion implicit in [undefj, Section 7] in two ways: the algebra $A$ need not have an algebraic identity, and the conclusion does not require the representation to be unital.
Proposition 4.3.
Let $A$ be a lattice-ordered algebra. Suppose there exist non-zero positive elements $e,p\in A_{+}$ satisfying:
  1. (i)
    $e^{2}=e$, $p^{2}=p$,
  2. (ii)
    $ep=pe=p$, and
  3. (iii)
    $p=\alpha e+x$, for some $x\in A$ with $x\wedge e=0$, and some $\alpha\in\mathbb{R}_{+}$.
If $\alpha\not\in\{0\}\cup\{1/n:n\in\mathbb{N}\}$, then $A$ has no faithful representation on $\mathcal{L}_{r}(X)$, for $X$ any Dedekind complete vector lattice.
Proof.
Suppose that $\Phi\colon A\to\mathcal{L}_{r}(X)$ is a faithful representation of $A$, where $X$ is a Dedekind complete vector lattice. Set $E=\Phi(e)$, $P=\Phi(p)$, and $T=\Phi(x)$. Since $\Phi$ is a lattice and algebra homomorphism, $E$ and $P$ are positive projections, and they are non-zero because $\Phi$ is faithful and $e,p\neq 0$. The relations
$ex=xe=x,$
which follow from $ep=pe=p$ together with $p=\alpha e+x$, give
$ET=TE=T.$
Applying $\Phi$ to (iii) yields $P=\alpha E+T$, and the disjointness $e\wedge x=0$ in $A$ transfers via $\Phi$ to $E\wedge T=0$ in $\mathcal{L}_{r}(X)$.
By 4.2, setting $Y=E(X)$, we obtain $\operatorname{id}_{Y}\wedge T|_{Y}=0$ in $\mathcal{L}_{r}(Y)$. Restricting the equality $P=\alpha E+T$ to $Y$ yields
$P|_{Y}=\alpha\operatorname{id}_{Y}+T|_{Y},$
where, by 4.1, $Y$ is a Dedekind complete vector lattice (non-zero, since $E\neq 0$), $P|_{Y}$ is a positive projection on it, and $T|_{Y}$ is disjoint from $\operatorname{id}_{Y}$. That is, $P|_{Y}$ has constant diagonal $\alpha$. By 3.1, $\alpha\in\{0\}\cup\{1/n:n\in\mathbb{N}\}$. ∎
Remark.
When $e$ is an algebraic identity and $\Phi$ is unital, 4.2 is not actually needed: $\Phi$ being unital then forces $E=\operatorname{id}_{X}$, so $Y=X$, and the operator $P=\Phi(p)$ already has constant diagonal $\alpha$ in $\mathcal{L}_{r}(X)$, allowing one to apply 3.1 directly. The role of 4.2 is precisely to handle the case in which $E=\Phi(e)$ is not the identity of $\mathcal{L}_{r}(X)$, which is what one encounters either when the algebra has no identity or when the representation is not required to be unital.
The following example, due to Wickstead [undefj, Section 7], exhibits Banach lattice algebras of dimension $2$ satisfying the conditions of 4.3. In its original form it answered [undefh, Question 5.3] and [undefi, Problem 2] in the case of unital representations; with the strengthened proposition above, the conclusion no longer depends on the representation being unital.
Example 4.4.
Let $1\leq\beta\leq 0$. Consider $\mathbb{R}^{2}$ with pointwise multiplication and positive cone
$P_{\beta}^{1}=\{\,(x,y)\in\mathbb{R}^{2}:0\leq x,\beta x\leq y\leq x\,\}.$
It is shown in [undefj, Theorem 3.1] that this is a lattice cone for which the product of positive elements is positive. Moreover, $(\mathbb{R}^{2},P_{\beta}^{1})$ can be endowed with a norm so as to make it a Banach lattice algebra [undefj, Proposition 8.3]. Also, the algebraic identity $e=(1,1)$ is positive.
Let $p=(1,0)$ and $x=(1,\beta)$. Note that $p^{2}=p$, $p\geq 0$, and that
$p=\frac{\beta}{\beta-1}e+\frac{1}{1-\beta}x.$
Let us check that $x\wedge e=0$. If $y=(y_{1},y_{2})$ and $x,e\geq y$, then
$\beta(1-y_{1})\leq\beta-y_{2}\leq 1-y_{1}\text{ and }\beta(1-y_{1})\leq 1-y_{2}\leq 1-y_{1},$
which give $\beta y_{1}\geq y_{2}\geq y_{1}$. Thus $y_{1}\leq 0$, which together with $\beta y_{1}\geq y_{2}\geq y_{1}$, implies that $y\leq 0$.
By 4.3, applied with $e=(1,1)$ and $p=(1,0)$, if
$\frac{\beta}{\beta-1}\not\in\{0\}\cup\{1/n:n\in\mathbb{N}\},$
that is, if $\beta\not\in\{0\}\cup\{-1/(n-1):n\in\mathbb{N},n\geq 2\}$, then $(\mathbb{R}^{2},P_{\beta}^{1})$ admits no faithful representation on any Dedekind complete vector lattice.
Acknowledgements
I would like to thank all the participants in the Workshop on Free Banach Lattices held at the University of Wuppertal in February 2026. One of the problem sessions at the workshop was devoted to this question. The enriching discussions not only provided valuable ideas, but also motivated me to further pursue this problem. I am particularly grateful to J. Glück for pointing out 2.3 and for organizing the workshop, and to P. Tradacete for his feedback on early versions of the paper. Finally, I would also like to thank Claude Opus 4.6 for providing the details of 4.2.